# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

### v2

@@ 5.1 Main Results
We conduct a parameter search experiment on the step $T_{max}$ of the alternating optimization in HLQ-GPTQ and investigate its impact on the final model performance. The experiments are carried out on LLaMA3.1-8B and Qwen3-8B, covering both w3g128 and w2g128 settings. The results are shown in the figure 8. We observe that increasing $T_{max}$ does not necessarily lead to continuous improvement. For both 3-bit and 2-bit quantization, setting $T_{max}$ = 10 yields a relatively good perplexity across both models, suggesting that a moderate optimization depth strikes a balance between convergence and stability. Specifically, for 3-bit quantization, the perplexity decreases steadily as $T_{max}$ increases up to 10 and then stabilizes, indicating that the alternating optimization approaches a convergence point. In contrast, the 2-bit setting exhibits a more irregular pattern, with notable oscillations in perplexity as $T_{max}$ increases. Perhaps incorporating certain regularization techniques could help the results converge further, which we leave for future work.
Figure 8: Effect of $T_{max}$ on perplexity under different quantization configurations. The figure shows the variation of perplexity with respect to $T_{max}$ for Llama3-8B and Qwen3-8B models under two quantization settings (W3G128 and W2G128).
5.2.3 Calibration Data Size for Efficient Finetuning.
We investigated the effect of the size of the calibration data on the final performance. As illustrated in Figure 9, our method yields substantial improvements as the calibration set size increases when it is below 1k. However, once the size exceeds 1k, the improvement becomes marginal. Since enlarging the calibration set introduces considerable memory overhead, especially during the block-wise stage, we set the calibration set size to 1k, thereby striking a balance between performance and efficiency.
Figure 9: Model performance of w2g128 Llama3.1-8b with different calibration size when finetuning.
5.2.4 Impact of Each Components in Efficient Finetuning.
Finetuning consists of two main stages: block-wise quantization and end-to-end tuning. We investigate the impact of each stage on the final model performance. The results are presented in Table 5.
Table 5: Impact of each component in Finetuning on Llama3.1-8B w2g128 quantization.
@@
5.3 Efficiency Evaluation.
5.3.1 Time and Memory Efficiency.
Table 6 reports the quantization time and peak GPU memory usage of HLQ-GPTQ and HLQ-Finetuning. HLQ-GPTQ is highly efficient, requiring only 0.15–0.40 hours for Llama2-7B and Llama2-13B, and 0.20 hours for Llama3.1-8B, with GPU memory usage between 7 and 10 GB. For Llama3.1-70B, it completes in 4 hours using 18 GB of memory, which fits easily within a single RTX 4090. HLQ-Finetuning also remains efficient. It uses 7–17 GB of memory for smaller models and 40 GB for Llama3.1-70B, and finishes in 1.5–4.5 hours for Llama2-7B and Llama2-13B, and 20 hours for Llama3.1-70B. Overall, both methods scale to large models while keeping memory usage low, making 2-bit and 3-bit quantization feasible on commodity GPUs.
Table 7 summarizes the quantization time and memory consumption of several 2-bit quantization pipelines on Llama3.1-70B during block reconstruction. Methods such as GPTQ and AWQ run relatively quickly but demand very large amounts of CPU memory. More advanced approaches, including OmniQ, Quip#, AQLM, DB-LLM, and EfficientQAT, require long processing times ranging from tens to hundreds of hours and extremely high CPU memory. Several of these methods also exceed the VRAM capacity of four RTX 4090 GPUs. In comparison, HLQ-GPTQ completes in 4 hours with 64 GB of CPU memory and 18 GB of GPU memory, offering a significantly more practical and resource-efficient solution. HLQ-Finetuning shows similar advantages: it keeps the same low memory requirements and completes within 20 hours, outperforming existing finetuning-based quantizers. These results demonstrate that HLQ is an efficient and scalable framework for large-scale 2-bit quantization under realistic hardware constraints.
Table 6: Quantization time and peak GPU memory usage of HLQ-GPTQ and HLQ-Finetuning.
Model HLQ-GPTQ HLQ-Finetuning
T Mem(3/2 bit) T Mem(3/2 bit)
llama2-7 0.15h 8/7GB 1.5h 9/7GB
llama2-13 0.40h 10/9GB 4.5h 17/13GB
llama3.1-8 0.20h 9/8GB 2h 10/8GB
llama3.1-70 4.00h 24/18GB 20h 40/32GB
Table 7: Comparison of quantization time and peak memory usage on W2g128 Llama3.1-70B across different quantization methods during block reconstruction. The results of Quip and Aqlm are inferred from LLaMA2-70B using 4k calibration data. >>96 indicates that the experiment could not be conducted on four RTX 4090 GPUs. The detailed experimental configurations for all methods can be found in Appendix D.
@@
GPTQ 2.50 180 28
AWQ 5 180 24
OmniQ 40 240 40
HLQ-GPTQ 4 64 18
Quip# 260 1200 >>96
Aqlm 300 950 >>96
DB-LLM 72 1200 >>96
EfficientQAT 50 300 30
HLQ-Finetuning 20 64 18
5.3.2 Data Efficiency.
Compared with previous approaches, our method demonstrates substantially higher data efficiency. For HLQ-GPTQ, we follow prior works and use only 128 calibration samples. For HLQ-Finetuning, although our final experiments adopt a calibration set size of 1k, even with only 128 samples our method still outperforms previous uniform quantization methods. Methods such as DB-LLM require up to 20k calibration samples, and Quip# and AQLM require 4k samples, whereas our approach achieves competitive performance with only 1k calibration samples, substantially reducing the need for calibration data.
@@ 5.4 Hardware Deployment on Edge Devices
5.4.1 Hardware Devices.
In this section, we primarily evaluate the inference performance of our framework on edge devices built on the ARM architecture. ARM-based processors are widely adopted in edge computing due to their highly integrated System-on-Chip (SoC) design and cache-friendly characteristics, which are advantageous for table-lookup–based operations. Moreover, ARM processors are well known for their low power consumption, making them a common choice in resource-constrained edge scenarios. For evaluation, we select two representative ARM-based chips: the Apple M2 from the Apple Silicon series and the RK3588 based on the ARM Cortex architecture. However, since HLQ can be converted into the BCQ format used in ShiftAddLLM (You et al., 2024) and AnyBCQ (Park et al., 2025a) through simple linear transformations, our quantization format is also compatible with GPU inference. We additionally provide experimental results on GPUs (see Appendix E.2 for details).
5.4.2 Evaluation Setup.
In terms of hardware evaluation, our experiments focus on the following aspects:
  • Numerical precision evaluation: Since our optimization algorithm is executed on the GPU, deploying the optimized model on edge devices may introduce numerical precision errors. Therefore, we evaluate the numerical accuracy of our framework. The experimental results demonstrate that the output of our framework achieves a cosine similarity greater than 99% compared to the GPU output, meeting the requirements for practical application.
  • Overhead of HLQ: we follow T-MAC’s design and adapt it to support HLQ. Experimental results show that the introduction of HLQ incurs no additional overhead to the Bit-serial LUT-based GEMM kernel.
  • End-to-End Speedup Comparison with AWQ: We compare our inference framework with AWQ for end-to-end speedup.
  • Compared with llama.cpp: We further compare our framework with the state-of-the-art edge-side inference framework, llama.cpp, to demonstrate the superiority of our method.
5.4.3 Numerical Precision Evaluation.
As mentioned in TensorRT (NVIDIA, 2023) and the MLPerf (Reddi et al., 2020), the cosine similarity between GPU and CPU outputs needs to be greater than 99% to meet the application standards. For the LLaMA2-7B model with 3-bit weight quantization, we select the output hidden states from the 9th, 18th, and 31st (final) layers on the M2 chip and compare them with the corresponding outputs on the GPU. The hidden states of the GPU denote $\mathbf{a}$, and the hidden states of the M2 denote $\mathbf{b}$. The cosine similarity is computed as
$\text{cos\_sim}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$
Results in Table 8 demonstrate that the cosine similarity between our framework and the GPU simulation exceeds 99.9%.
Table 8: Cosine similarity(%) between simulated quantization on GPU and actual quantization on the M2 chip at the 9th, 18th, and 31st(final) layers of the LLaMA2-7B model with 3-bit weight quantization.
Model Prompt L9 L18 L31
LLaMA2-7B-3Bit 128 99.9994 99.9990 99.9982
256 99.9984 99.9986 99.9970
512 99.9988 99.9978 99.9960
5.4.4 Overhead of HLQ.
Table 9: Comparison of GEMM latency (ms) between uniform quantization and HLQ formats using the T-MAC kernel.
Chip Kernel $11008\times 4096$ $4096\times 32000$
T-MAC-HLQ 5.72(± 0.11)5.72\,($\pm\,$0.11) 22.23(± 0.16)22.23\,($\pm\,$0.16)
T-MAC-HLQ 1.86(± 0.04)1.86\,($\pm\,$0.04) 4.49(± 0.05)4.49\,($\pm\,$0.05)
Method W G Mem(GB) Speedup
[…]
Table 11: Comparison of prefill and decode throughput (tokens/s) between ELUTQ and llama.cpp on the Apple M2 chip. All experiments are conducted with 4 threads and a batch size of 1. More results can be found in Appendix E.1
Model Framework Wbits BPW Input,Output Prefill Decode
llama2-7B llama.cpp Q3_K_M 3.91 128,128 19.87 15.65
256,128 19.75 15.53
Q3_K_S 3.50 128,128 17.60 15.44
256,128 17.43 15.00
Q2_K_S 2.50 128,128 14.82 13.35
256,128 14.54 13.16
ELUTQ W3(g128) 3.50 128,128 20.60 16.63
256,128 21.54 16.21
W2(g128) 2.37 128,128 25.06 20.48
256,128 25.94 20.25
5.4.5 End-to-End Speedup Comparison with AWQ
Here, we compare ELUTQ with the uniform-quantization-based inference framework AWQ in terms of end-to-end speedup on the LLaMA3.1-8B model. All experiments are conducted on the Apple M2 chip. As shown in Table 10, under the 3-bit setting, AWQ achieves a 2.0× speedup over FP16, while ELUTQ achieves 2.5×. Under the 2-bit setting, AWQ achieves 1.6×, whereas ELUTQ reaches 3.4×. These results indicate that AWQ suffers from significant dequantization overhead at lower bit widths, resulting in reduced speedup at 2 bits, whereas our framework continues to improve efficiency as the bit width decreases.
5.4.6 Compared with llama.cpp.
Directly comparing inference speed with llama.cpp is not entirely fair, since llama.cpp adopts its own quantization format, GGUF, which is incompatible with our quantization scheme. To ensure a fair comparison, we instead use bits per weight (BPW) to measure model size, focusing on how inference speed changes as the model size varies.
Table 11 presents the results on LLaMA2-7B and Qwen2-1.5B. We observe that under lower-bit settings, llama.cpp becomes slower as the model size decreases. For example, Q3_K_S is slower than Q3_K_M and Q2_K_S is slower than Q3_K_S. This phenomenon mainly arises from the substantial decoding overhead associated with low-bit quantization. In contrast, our framework benefits from the Bit-serial LUT-based GEMM paradigm, achieving nearly linear improvements in inference speed as the average bit-width decreases. Specifically, when BPW = 3.5, ELUTQ outperforms llama.cpp by approximately 20% in the prefill stage and 10% in the decode stage. When BPW $\approx$ 2.5, the speedup increases to about 45% and 25%, respectively.
Moreover, the improvement is more pronounced in the prefill stage than in the decode stage. This is because the prefill stage is computation-intensive and primarily involves GEMM operations, while the decode stage is memory-intensive and mainly consists of GEMV operations. In the prefill stage, our method replaces multiplication in matrix multiplication with efficient lookup and sum operations, significantly accelerating computation. In contrast, the decode stage is memory-intensive, where runtime is dominated by memory access. Since the Bit-serial LUT-based GEMM requires additional table lookups, its memory access cost is slightly higher than that of traditional GEMV, leading to smaller speed gains in the decode stage.
6 Limitations and Discussion
@@
  • Weight-only quantization. The current study focuses exclusively on weight only quantization. Extending our framework to weight activation quantization could further enhance inference efficiency and reduce memory consumption. This remains an open and valuable direction for future research.
  • Integration with other quantization frameworks. At present, HLQ has been integrated only with a limited set of efficient quantization methods. Combining HLQ with Quantization-aware Training or LoRA-based adaptation may lead to additional gains in quantization accuracy, which we plan to explore in subsequent work.
@@ 7 Conclusion
In this paper, we propose ELUTQ, an efficient quantization framework designed for deploying large language models on edge devices. The framework is carefully aligned with the computation pipeline of Bit-serial LUT-based GEMM and introduces a novel Hierarchical Linear Quantization method that better captures the weight distribution compared to traditional uniform quantization. Moreover, HLQ can be seamlessly integrated with existing quantization techniques, including post-training and fine-tuning, to further enhance model accuracy. To reduce both GPU and CPU memory consumption, we further optimize the model and data loading pipeline in the quantization process, enabling quantization algorithms to operate efficiently under limited computational resources. Finally, we develop a pure C++ inference framework that enables accurate and efficient on-device inference, facilitating the practical deployment of quantized models in edge scenarios.
@@ Appendix A Gradient-Based Search for Hierarchical Linear Quantization
To ensure training stability, we enforce a non-negativity constraint on the scaling factors and apply dynamic range clipping to the zero-point values, which effectively prevents training divergence. From an optimization perspective, this approach is conceptually aligned with the straight-through estimator (STE) (Liu et al., 2022) philosophy. However, a key distinction lies in the fact that instead of relying on handcrafted gradient approximation rules, we carefully design the quantization expression itself to allow gradients to flow naturally to the continuous parameters, resulting in a more elegant approximation of gradient flow.
@@
We conduct an experiment where HLQ-GPTQ is implemented using both alternating optimization and gradient-based search methods. As shown in Table 12, both methods have similar results. Therefore, considering algorithmic robustness, we recommend adopting the alternating optimization method in practical applications. Besides, the alternating optimization approach can complete model quantization within a very short time (Table 13).
Table 12: Performance comparison of HLQ-GPTQ using alternating optimization and gradient-based optimization across different models and bit configurations. perplexity($\downarrow$) is reported.
Method # W # G BPW LLaMA2-7 LLaMA2-13 LLaMA3.1-8 Qwen3-8
wikitext2 c4 wikitext2 c4 wikitext2 c4 wikitext2 c4
Baseline 16 - 16 5.47 6.97 4.88 6.47 6.15 8.89 9.72 13.30
Alternating Optimization 2 128 2.37 9.95 15.02 7.12 9.92 16.93 25.44 18.15 24.63
Gradient-based Method 2 128 2.37 9.90 14.89 7.02 9.75 17.32 27.14 18.13 24.60
Alternating Optimization 3 128 3.5 5.97 7.74 5.14 6.91 7.24 10.80 10.45 14.15
Gradient-based Method 3 128 3.5 5.97 7.74 5.14 6.90 7.41 10.85 10.53 14.22
Table 13: Runtime (h) comparison of GPTQ, HLQ-GPTQ (Alternate), and HLQ-GPTQ (Gradient).
Model GPTQ HLQ-GPTQ(Alternate) HLQ-GPTQ(Gradient)
LLaMA2-7B 0.20 0.15 0.20
LLaMA2-13B 0.40 0.35 0.40
LLaMA3.1-8B 0.25 0.20 0.25

### v3

@@ 5.2.2 Hyparameter Study for $T_{max}$
We conduct a hyparameter search experiment on the step $T_{max}$ of the alternating optimization and investigate its impact on the final model performance. The experiments are carried out on LLaMA3.1-8B and Qwen3-8B, covering both W3G128 and W2G128 settings. The results are shown in the figure 6. We observe that increasing $T_{max}$ does not necessarily lead to continuous improvement. For both 3-bit and 2-bit quantization, setting $T_{max}$ = 10 yields a relatively good perplexity across both models.
Figure 6: Impact of the alternating optimization step $T_{max}$ on quantization performance. Results are reported in terms of perplexity ($\downarrow$) on LLaMA3.1-8B and Qwen3-8B under W3G128 and W2G128 settings.
5.2.3 Impact of Each Stage in Efficient Finetuning
Finetuning consists of two stages: block-wise quantization and end-to-end finetuning. We evaluate their individual and combined effects in Table 4. Without either stage, the quantized model exhibits severe degradation. Block-wise quantization alone substantially improves both perplexity and accuracy, whereas end-to-end finetuning alone provides limited benefits. Combining both stages achieves the best performance.
@@ 5.2.3 Impact of Each Stage in Efficient Finetuning
Table 4: Impact of each part in finetuning on LLaMA3.1-8B W2G128 quantization.
5.3 Efficiency Evaluation
Table 5 summarizes the quantization time and memory consumption of several 2-bit quantization pipelines on LLaMA3.1-70B. While methods such as GPTQ and AWQ complete quantization relatively quickly, they require large amounts of CPU memory. Some finetuning approaches, including OmniQ, Quip#, AQLM, DB-LLM, and EfficientQAT, incur substantially longer processing times ranging from tens to hundreds of hours and demand extremely high CPU memory. In comparison, ELUTQ optimizes only a small number of parameters during fine-tuning under an efficient alternating optimization algorithm and a two-stage strategy, and by incorporating lazy loading and disk offloading into the quantization pipeline, it maintains low memory requirements and completes quantization within 40 hours. We further provide a parameter efficiency evaluation in Appendix G.
@@ 5.3 Efficiency Evaluation
GPTQ 1.50 180 32
AWQ 2.0 180 24
OmniQ 50 240 48
Quip# 280 1200 320
Aqlm 320 950 320
DB-LLM 120 1200 320
EfficientQAT 70 400 48
ELUTQ 40 64 48
Table 5: Comparison of quantization time and peak memory usage on W2G128 Llama3.1-70B across different quantization methods. All experiments are conducted on RTX4090s. More details can be found in Appendix G.
@@ 5.4 Hardware Deployment on Edge Devices
Our evaluation includes consumer-grade CPU chips such as Rk3588 and Apple M2, as well as consumer-grade GPUs including the RTX 3090 and RTX 4090. In addition, we conduct experiments on server-class accelerators, including the A100 and H100. Full results can be found in Appendix H.
As shown in figures 7 and 8, we compare our method against llama.cpp for CPU inference and three representative GPU baselines, including AWQ with uniform quantization, SqueezeLLM with LUT-based dequantization, and Quip# with codebook-based quantization, using FP16 as the reference. On the CPUs, our 2-bit LLaMA3.1-8B model achieves an approximately 1.4× speedup over llama.cpp. On the GPUs, we observe that Quip# performs worse than FP16 due to highly irregular memory accesses caused by codebook lookups, while SqueezeLLM, despite leveraging LUT-based dequantization, still falls behind AWQ. However, AWQ incurs increased weight unpacking overhead at lower bit-widths, resulting in inferior performance at 2-bit compared to 3-bit. In contrast, ELUTQ, built upon the bit-serial LUT-based GEMM paradigm, consistently achieves higher throughput than dequant-based methods, with inference speed further improving as the bit-width decreases. Notably, the performance gains of ELUTQ become more pronounced for larger models and lower bit-widths.
Figure 7: End to End Evaluation (tokens/s) of generating a sequence lengeth of 128 on Apple M2 and RK3588 using 4 threads. Experiments were conducted with a batch size of 1.
Figure 8: End to End Evaluation (tokens/s) of generating a sequence lengeth of 1024 on RTX3090, RTX4090, A100, H100. Experiments were conducted with a batch size of 1. FP16 inference of LLaMA3.1-70B is not feasible on A100 and H100 due to memory constraints.
6 Limitation and Discussion
@@
  • Weight-only quantization. The current study focuses exclusively on weight only quantization. Extending our framework to weight-activation quantization could further enhance inference efficiency and reduce memory consumption.
  • Other quantization frameworks. Combining HLQ with Quantization-aware Training or LoRA-based adaptation may lead to additional gains in quantization accuracy, which we plan to explore in subsequent work.
@@ 7 Conclusion
In this paper, we propose ELUTQ, an efficient quantization framework designed for deploying large language models on edge devices. The framework is carefully aligned with the computation pipeline of Bit-serial LUT-based GEMM and introduces a novel Hierarchical Linear Quantization method that better captures the weight distribution compared to traditional uniform quantization. Moreover, we designed an optimized quantization pipeline for HLQ to enhance model accuracy and reduce both GPU and CPU memory consumption during the quantization process. Finally, we support end-to-end inference to facilitate the practical deployment of quantized models in edge scenarios.
@@ Acknowledgments
Appendix
Appendix Overview
Appendix A:Gradient-Based Search for Hierarchical Linear Quantization
We introduce an alternative gradient-based approach for searching the scale and zero-point of HLQ.
Appendix B: Detailed Computational procedure for Bit-Serial Lut-Based GEMM
We provide a detailed description of the computation procedure for Bit-Serial LUT-based GEMM.
Appendix C: System-level Optimizations for the Quantization Pipeline
We present key technical components, including lazy loading, chunked computation, and disk offloading, to enable model quantization for large-scale models under limited CPU and GPU memory budgets.
Appendix D: Design of High Performance CPU and GPU kernels
Core design details of the GPU and CPU kernels.
Appendix E: Compression Ratios and Actual Memory Footprint of Quantized Models
Details of Compression Ratios and Actual Memory Footprint.
Appendix F: Compared with Non-Uniform and Quantization-Aware Training Methods
Further comparison of ELUTQ with codebook-based non-uniform quantization methods and QAT approaches.
Appendix G: More details of Efficiency Evaluation
Dtailed experimental configurations used to measure the computational memory consumption.
Appendix H: Full Results of Inference Performance on Edge Devices
We comprehensively present the inference performance of ELUTQ on both CPU and GPU, including kernel-level evaluation and end-to-end evaluation.
Appendix I: Full Results on Model Accuracy
We present full results on model accuracy.
@@ Appendix A Gradient-Based Search for Hierarchical Linear Quantization
To ensure training stability, we enforce a non-negativity constraint on the scaling factors and apply dynamic range clipping to the zero-point values, which effectively prevents training divergence. From an optimization perspective, this approach is conceptually aligned with the straight-through estimator (STE) Liu et al. [2022] philosophy. However, a key distinction lies in the fact that instead of relying on handcrafted gradient approximation rules, we carefully design the quantization expression itself to allow gradients to flow naturally to the continuous parameters, resulting in a more elegant approximation of gradient flow.
@@
Appendix B Detailed Computational procedure for Bit-Serial Lut-Based GEMM
Here, we provide a detail computational procedure for Bit-serial Lut-based GEMM. There are two primary computational paradigms for LUT-based GEMM, as illustrated in the Figure 9.
Figure 10: Detail pipeline of Bit-serial Lut-based GEMM.
Figure 9(c) illustrates another GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the scheme in Figure 2(b), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. The detail computational procedure is shown in Figure 10. First, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{W_{0},W_{1},...,W_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. For example, for integer values (9, 7, 6, 3) with binary representations (1001, 0111, 0110, 0011), the matrix for the lowest bit is (1, 1, 0, 1), and the matrix for the highest bit is (1, 0, 0, 0). This decomposition is performed offline, incurring no runtime overhead. During inference, for an activation vector of the group size $g$, the system precomputes the dot products between this activation vector and all $2^{g}$ possible combinations of single-bit weights, storing the results in the LUT. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation. This paradigm has been demonstrated to offer high computational efficiency and energy efficiency Park et al. [2025b]; Wei et al. [2025]. For example, FIGLUT Park et al. [2025b] optimized the table structure for GPU architectures to avoid bank conflicts, while T-MAC leveraged CPU vectorized lookup instructions (AVX2/NEON) to enable efficient LUT operations on CPUs.
Appendix C System-level Optimizations for the Quantization Pipeline
C.1 Lazy Loading
In both GPTQ’s Hessian-based quantization and Efficient Finetuning’s block reconstruction, the quantization unit is a Transformer block. Quantizing a single block does not depend on other blocks, neither the already quantized nor the unquantized ones. Therefore, instead of loading the entire model into memory, we can load only one transformer block at a time from disk. Thanks to the sharded weight storage mechanism in the Transformers Wolf et al. [2020] framework, we can leverage the weight mapping file (e.g., model.safetensors.index.json) to selectively load a specific block from disk for quantization. Once the quantization of a block on the GPU is completed, it is immediately offloaded to the CPU to reduce GPU memory usage.
C.2 Chunked Computation of HLQ Parameters
When computing the initial quantization parameters of HLQ for a weight matrix, the optimization within each weight group is independent. For extremely large weight matrices, we can split the matrix into smaller chunks and perform parameter search independently for each chunk. As shown in the figure, for the largest matrix in Llama3.1-70B, the down_proj of shape [28672, 8192] performing HLQ parameter search without batching leads to a peak memory usage of approximately 30 GB for 2-bit quantization and 35 GB for 3-bit quantization. In contrast, when the search is executed in two chunks, the peak memory drops to about 22 GB for 2-bit and 24 GB for 3-bit quantization. This chunked computation strategy significantly reduces GPU memory consumption and achieves approximately linear memory scaling with respect to partition granularity.
C.3 Offloading Hidden States to Disk
C.4 Overhead of Disk Access
Using lazy loading and loading training data directly from disk requires disk access during the execution of quantization algorithms. This inevitably introduces additional memory access overhead. Even with NVMe-based communication protocols, disk bandwidth remains far below that of the CPU. As shown in Table 6, we evaluate ELUTQ on LLaMA3.1-8B, Qwen3-32B, and LLaMA3.1-70B. The impact of disk access is relatively minor compared to the overall tuning time. Specifically, the tuning time increases by 25% for LLaMA3.1-8B, 12% for Qwen3-32B, and only about 5% for LLaMA3.1-70B. Therefore, when performing ELUTQ on very large models, the overhead caused by disk access can be considered negligible relative to the total tuning time.
Table 6: Comparison of tuning times with and without disk usage. All models are tuned under the same quantization configuration w2g128, using 1k calibration dataset.
Method Model disk(✗) disk(✓) Overhead (%)
ELUTQ llama3.1-8B 2.00h 2.50h 25%
Qwen3-32B 8.25h 9.30h 12%
llama3.1-70B 19.00 20.00h 5%
Appendix D Design of High Performance CPU and GPU kernels
D.1 Weight-Rearrange.
D.2 Mirror Storage.
Given the formulation:
$\hat{\mathrm{W}}=\sum_{j=0}^{q-1}s_{j}\cdot b_{j}+z,b_{j}\in\{0,1\}^{n},$ (10)
D.3 Table Quantization.